How to write this on Excel?

1.99-2.0 0.05 /25 Z=-1.0
Answer #1

Ans:

z=\dfrac{\bar{x}-\mu }{\sigma /\sqrt{n}}

Use below to write in excel:

z=(1.99-2)/(0.05/SQRT(25))

z=-1

P(z<-1)=normsdist(-1)=0.1587